OpenStudy (anonymous):

How do i do these? Evaluate: 1. cos 87 degree 20' 2. sin 6 degree 50' 3. sec 51 degree 30' 4. csc 62 degree 10' 5. cos 24 degree 43'

OpenStudy (anonymous):

with a calculator but one step first

OpenStudy (anonymous):

Or wolfram

OpenStudy (anonymous):

but I don't trust him -.-

OpenStudy (anonymous):

\[20'=\frac{20}{60}=\frac{1}{3}=0.333333...\]so you need to put in \(\cos(87.3333)\) because your calculator works with decimals, not minutes

OpenStudy (anonymous):

or try this http://www.wolframalpha.com/input/?i=cos%2887^o20%27%29